How do I solve this -x-y=-1 & -x-2y=-10
Answer by Jk22(389)

-x-y=-1
-x-2y=-10


(1)-(2) : y = -1-(-10) = 10 - 1 = 9, and from equation 1 : x = 1-y = -8

In equation 2 : 8-18 = -10